OpenStudy (anonymous):

Solve:

OpenStudy (anonymous):

\[2^{2x}-2^{x+1}-15=0\]

OpenStudy (anonymous):

@ganeshie8 @triciaal

OpenStudy (anonymous):

@Hero @Kainui

OpenStudy (triciaal):

one approach put the number to the right take the logs of each side

OpenStudy (anonymous):

what number goes to the right?

OpenStudy (triciaal):

when you subtract logs you divide the numbers

OpenStudy (anonymous):

I know but you said to put the number on the right. What number is that? the 2^2x? 2^x-1?15?

OpenStudy (triciaal):

constant 15

OpenStudy (triciaal):

isolating the x

OpenStudy (triciaal):

using base 2

OpenStudy (triciaal):

@Nnesha can you help?

OpenStudy (anonymous):

sorry can u draw it out? I know its a lot to ask for

OpenStudy (triciaal):

@ganeshie8 please help

OpenStudy (triciaal):

@UnkleRhaukus please help

OpenStudy (unklerhaukus):

\[2^{2x}-2^{x+1}-15=0\\ 2^{2x}-2^{x+1}=15\\2^x(2^x-2)=15\]

OpenStudy (anonymous):

then what?

OpenStudy (unklerhaukus):

i'm not sure i got stuck

Nnesha (nnesha):

need wrong answer ?? tell me so i can try :(

OpenStudy (anonymous):

Tell you what?

Nnesha (nnesha):

nvm nothing :)

Nnesha (nnesha):

idk how u got 2^x(2^x -2) ??

Nnesha (nnesha):

you should move  this -2 to other side then its become positive then you can take log